The vapor pressure of a substance is 760.0 torr at 115°C and 14.6 torr at 45°C. What is ΔHvap of the substance?
Answer #1

= 388 K According to Clausins-clapeysson equation lul- P2 Alvap T;=6273+115) K Pi R 7 lu 14.6 Alvap T, =(273+ 45) K 760 Atvap